A man weighs 0.70kN on Earth. What is his mass in kilograms?
1
Expert's answer
2018-06-05T08:46:08-0400
We can find the mass of the man in kilograms from the definition of the weight:
W=mg,
here, W is the weight of the man, m is the mass of the man, g=9.8 m⁄s^2 is the acceleration due to gravity.
Then, we get:
m=W/g=(0.7∙〖10〗^3 N)/(9.8 m/s^2 )=71 kg.